OpenStudy (anonymous):

lim x → π/2 6 cos x cot x

myininaya (myininaya):

first write cot(x) in terms of sin and cos

OpenStudy (anonymous):

aha

myininaya (myininaya):

then just plug in pi/2

myininaya (myininaya):

since the function is continuous at pi/2

OpenStudy (anonymous):

ok thanks

myininaya (myininaya):

what did you get?

OpenStudy (anonymous):

6

myininaya (myininaya):

hmm...

myininaya (myininaya):

what is cos(0)?

myininaya (myininaya):

oops i mean cos(pi/2)?

OpenStudy (anonymous):

lim x → 0 sin8 x x

myininaya (myininaya):

Use your unit circle if you have to

myininaya (myininaya):

@althary we are still on the first one 6 is not a correct answer

OpenStudy (ikram002p):

no ne2eded to write is as sin and cos function , cot is continues at pi/ thus the lim x → π/2 6 cos x cot x = cos π/2 cot π/2

myininaya (myininaya):

what is cos(pi/2)?

OpenStudy (anonymous):

it's correct, I'm sure

myininaya (myininaya):

Well I'm telling 6 is not correct because cos(pi/2)=0

OpenStudy (anonymous):

u r gona cancle out cos with cosi and then u will end up with sin pi/2 whih is 1 and then times 6 so....

myininaya (myininaya):

you have 6 cos(x)*cos(x)/sin(x) nothing cancels

myininaya (myininaya):

The only reason I had you write it in terms of sin and cos is to make it easier for you to determine that it was actually continuous at pi/2 since sin(pi/2)=1

OpenStudy (anonymous):

read the q again

OpenStudy (unklerhaukus):

cot = cos/ sin, @alalthary, not sin/cos

myininaya (myininaya):

you might want to read what you type if you think i'm reading it wrong

myininaya (myininaya):

did you mean 6cos(x)/cot(x)?

OpenStudy (anonymous):

yes exactly

myininaya (myininaya):

well everyone has been reading 6cos(x)cot(x) :p

OpenStudy (anonymous):

ooh sorry then

myininaya (myininaya):

in any case \[6 \cos(x) \cdot \frac{\sin(x)}{\cos(x)} \] you would then be right

myininaya (myininaya):

the cosine's will cancel

myininaya (myininaya):

and sin(pi/2)=1 and 6*1=6

OpenStudy (anonymous):

yes mam..

myininaya (myininaya):

ok so your next question did you say sin(8x)/x as x goes to 0?

OpenStudy (anonymous):

yeah sin^*8/x

myininaya (myininaya):

sin is to the eight power? let me write in latex

myininaya (myininaya):

\[\lim_{x \rightarrow 0}\frac{\sin(8x)}{x} \text{ or } \lim_{x \rightarrow 0}\frac{\sin^8(x)}{x} ?\] 1st or second?

myininaya (myininaya):

or neither can be choice

myininaya (myininaya):

in any of those two cases you will need to recall the following limit \[\lim_{u \rightarrow 0} \frac{\sin(u)}{u}=1\]

OpenStudy (anonymous):

the second one yeah

OpenStudy (anonymous):

sin to the power 8

OpenStudy (anonymous):

what's up, what to do with powers??

myininaya (myininaya):

so if u=8x we still have x going to 0 since u=8x as u->0 implies 0=8x which means x=0 which means x->0 \[\lim_{8x \rightarrow 0}\frac{\sin(8x)}{8x} =1 \text{ or you could just say } \lim_{x \rightarrow 0} \frac{\sin(8x)}{8x}=1\] oh well I will show you how to do the first one and you may be able to do the second one we will see \[\lim_{x \rightarrow 0}\frac{\sin(8x)}{x} =\lim_{x \rightarrow 0} 8 \frac{\sin(8x)}{8x}=8(1)=8\] But we have the second one: \[\lim_{x \rightarrow 0}\frac{\sin^8(x)}{x}\] I will rewrite this giving you a major hint: \[\lim_{x \rightarrow 0}\frac{(\sin(x))^8}{x^8} \cdot x^7 \]

myininaya (myininaya):

Tell me if you see where my hint is going or not

myininaya (myininaya):

that thing at the bottom of that big ugly post

OpenStudy (anonymous):

didnt undrstand

myininaya (myininaya):

\[\text{ Let } n \neq 0 ; \lim_{x \rightarrow 0} (\frac{\sin(x)}{x})^n =(\lim_{x \rightarrow 0}\frac{\sin(x)}{x})^n=1^n=1\]

myininaya (myininaya):

not that is pretty so I seen sin^8(x)/x looks kinda similar to that so I tried to use that form in it the only way I could do that is if I wrote 1/x as x^7/x^8

myininaya (myininaya):

\[\lim_{x \rightarrow 0}(\frac{\sin(x)}{x})^8 \cdot x^7 \]

OpenStudy (anonymous):

i didn't understand

myininaya (myininaya):

which part?

myininaya (myininaya):

Do you know that sin(x)/x goes to 1 as x goes to 0?

OpenStudy (anonymous):

yes i know but whats up with the power 8 !!!

myininaya (myininaya):

The problem has a power 8 in it: \[\lim_{x \rightarrow 0}\frac{\sin^8(x)}{x^8} =1^8=1 \]

myininaya (myininaya):

\[\lim_{x \rightarrow 0} \frac{\sin^8(x)}{x}\] This is your problem to use what I just said I need to multiply x^7 on both top and bottom

myininaya (myininaya):

\[\lim_{x \rightarrow 0}\frac{\sin^8(x)}{x} \cdot \frac{x^7}{x^7} =\lim_{x \rightarrow 0}\frac{x^7 \sin^8(x)}{x^8}\]

myininaya (myininaya):

\[\lim_{x \rightarrow 0}x^ 7 \cdot ( \lim_{x \rightarrow 0} \frac{\sin(x)}{x})^8 \]

OpenStudy (anonymous):

0 right

myininaya (myininaya):

yes 0*1=0